OpenStudy (klimenkov):

\[\int_{-\frac{\pi}2}^{\frac{\pi}2} \frac{\sin^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx\]

ganeshie8 (ganeshie8):

even/odd stuff ha ?

OpenStudy (anonymous):

2014 is order of derivative ?

hartnn (hartnn):

\(\Large \int_a^b f(x) dx= \int_a^bf(a+b-x)dx \) use this!

OpenStudy (klimenkov):

@BSwan 2014 is the power.

hartnn (hartnn):

after you have proved that your function is even

OpenStudy (anonymous):

aha cool :)

hartnn (hartnn):

\(\Large \int_{-a}^a f(x) dx = 2\int_0^a f(x)dx\) if f(x) is even function

hartnn (hartnn):

did u get what to do ?

OpenStudy (klimenkov):

\[2\int\limits_0^{\frac{\pi}2} \frac{\sin^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx\]What's next?

hartnn (hartnn):

\(\Large \int_a^b f(x) dx= \int_a^bf(a+b-x)dx\) replace x by pi/2 - x

hartnn (hartnn):

\(I = 2\int\limits_0^{\frac{\pi}2} \frac{\sin^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx ... ... (A)\) just giving a label, to be used later

OpenStudy (klimenkov):

\[2 \int\limits_0^{\frac{\pi}2} \frac{\cos^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx\]

OpenStudy (klimenkov):

Oops...

hartnn (hartnn):

\(I = 2 \int\limits_0^{\frac{\pi}2} \frac{\cos^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx ... ... (B)\) Add (A) and (B)

OpenStudy (klimenkov):

\[I = -2 \int\limits_0^{\frac{\pi}2} \frac{\cos^{2014}x}{\sin^{2014}x + \cos^{2014}x} \, dx\]

hartnn (hartnn):

what ? why negative ?

hartnn (hartnn):

dx remains as dx

hartnn (hartnn):

i am NOT doing any substitution

OpenStudy (klimenkov):

Yeah, everything is okay. My fault. \[I + I = 2\int_0^{\frac\pi2}dx,\]\[I=\frac\pi2.\]Very nice, thank you.

hartnn (hartnn):

welcome ^_^